News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 17 November 2000 5.00pm EST

During last week the observing schedule was resumed on Nov 11 following the solar event of Nov 8. The schedule for the coming week was then re-planned to include an observation of the X-ray transient Aql X-1 which was triggered as a Target Of Opportunity (TOO).

A real-time test with ACIS was conducted Nov 13 to demonstrate a more efficient approach to downlinking raw frame data. Assessment of the results indicate that the test was successful. The approach will be used in assessing further squeegee mode test data. Other real-time activities included uplink of a non-checksum software patch on Nov 16 to clear the eclipse counter that incremented during the lunar eclipse of Oct 27 and re-enable SCS 12, a routine ephemeris update, and a check of the CPE warm start count that showed the value remained at 1.

Preparations are complete for the 2000 Leonids meteor shower passage expected to peak for Chandra at ~12:45am EST on 18 November. The predicted Zenith Hourly Rate this year is ~700 compared with 3700 observed during the 1999 passage. In order to minimize possible exposure to particles, Chandra is scheduled to observe targets within 7 degrees of the anti-radiant for the duration of the peak.

The replanned schedule including the TOO and the observations for the coming week are as follows. The observation of PMN J1131+0455 was impacted by the TOO and will be rescheduled in an upcoming week.

Aql X-1 (TOO)           HRC-S/LETG   Nov 18
Eta Car                 ACIS-S/HETG
NGC 3368                ACIS-S       Nov 20 
Radiation Belt
3C 66B                  ACIS-S
HDF North               ACIS-I       Nov 21 
Radiation Belt                       Nov 23 
HDF (Quick Look)        ACIS-I       Nov 24 
Cen X-3                 ACIS-I       Nov 25 
Crab Nebula             ACIS-S
Radiation Belt                       Nov 26 
KI 284                  ACIS-S
M31 Transient           ACIS-S


All spacecraft subsystems continued to operate nominally.

Return to Status Reports